DogGirl27:

A rectangle’s width is 2 meters shorter than its length, l. Its area is 168 square meters. Which equation can be used to find the length of the rectangle? l(l – 2) = 168 l(l + 2) = 168 2l 2 = 168 l 2 = 168

Alex1513:

Let w = width Let l = length Let A = area w = l - 2 A = l * w also, A = 168 equating the two values of A l * w = 168 l * (l - 2) = 168 l^2 - 2l = 168 l(l - 2) = 168